In the compression stroke of an internal combustion engine the heat rejected to

the cooling water is 45 KJ/Kg and the work input is 90 KJ/Kg. Calculate the

Change in internal energy of the working fluid stating whether it is a gain or

loss


1
Expert's answer
2021-02-05T14:06:25-0500

According to the first law of thermodynamics


"Q=\Delta U+W\to \Delta U=Q-W=-45-(-90)=45\ (kJ/kg)"


It is a gain.